OpenStudy (anonymous):

Harry can rake the leaves in the yard 8 hours faster than his little brother Jimmy can. If they work together, they can complete the job in 3 hours. Using complete sentences, explain each step in figuring out how to determine the time it would take Jimmy to complete the job on his own.

OpenStudy (anonymous):

Solve the following for jimmy:\[\frac{1}{\text{jimmy}}+\frac{1}{\text{jimmy}-8}=\frac{1}{3} \]There are two solutions. Only one is valid.

OpenStudy (anonymous):

1/x+1(x+8)=1/3 so harry would be x+8

OpenStudy (anonymous):

1 job in jimmy hours plus 1 job in jimmy minus 8 hours has to be equal to 1 job in 3 hours.

OpenStudy (anonymous):

1/x+1/(x-8)=1/3 You left out the division sign in the second term and it's x-8 in the divisor. x = the time that jimmy takes to complete job by himself. Harry takes 8 hours less to do the job my himself, or, x-8 .

OpenStudy (anonymous):

It is not always necessary to express formulas in terms of x and y just because that is done in today's classroom.